OpenStudy (anonymous):

Correct ? Factor by grouping: xy - 2y + 2x - 4. (x - 2)(y + 2)

Directrix (directrix):

xy - 2y + 2x - 4 = y(x-2) + 2(x-2) = (y +2)(x-2) = (x-2)(y+2) Yes, your answer is correct.